\begin{thebibliography}{99}

\bibitem{bvg}Bini, G., van Geemen, B., Tyler, K., L., Mirror quintics, discrete symmetries and Shioda maps, e-print math.AG/0809.1791

\bibitem{bh}Berglund, P., H\"{u}bsch, T., A generalized construction of mirror manifolds. {\em Essays on mirror manifolds}, editor S-T Yau, Int. Press, Hong Kong 1992, 388--407.

\bibitem{cdk} Candelas, P., De la Ossa, X., Katz, S., Mirror symmetry for Calabi-Yau Hypersurfaces in Weighted
 ${\mathbb P}_4$ and extensions of Landau-Ginzburg Theory, {\em Nucl. Phys B} {\bf 450} (1995), no. 1-2, 267--290.

\bibitem{cls} Candelas, P., Lynker, M., Schimmrick, R., Calabi-Yau manifolds in weighted ${\mathbb P}_4$, {\em Nuclear Phys.} B
{\bf 341} (1990), no. 2, 383--402.

\bibitem{CK}Cox, D., A., Katz, S., Mirror symmetry and algebraic geometry, Mathematical surveys and monographs {\bf 68}, AMS, Providence, 1999.

\bibitem{Di} Dimca, A., Singularities and coverings of weighted complete intersections, {\em J. reine u. angew. Math.} {\bf 366} (1986) 184--193.


\bibitem{Dol} Dolgachev, I., Weighted projective space, in {\em Group Actions and vector fields} (J. B. Carrel ed.), Lecture Notes in Mathematics {\bf 956} , Springer-Verlag, New York 1982, 34--71.

\bibitem{dgj} Doran, C., Greene, B., Judes, S., Families of quintic
Calabi-Yau 3-folds with discrete symmetries, e-print hep-th/0701206.

\bibitem{GKZ} Gelfand, I., Kapranov, M., and Zelevinsky, A., Generalized Euler integrals and A-hypergeometric functions, {\em Adv in Math.} {\bf 84} (1990), 255-271.

\bibitem{gp} Greene, P., Plesser, M.R., Duality in Calabi-Yau Moduli space, {\em Nuclear Phys.} B {\bf 338} (1990), no. 1, 15-37.

\bibitem{gpr}Greene, B.R., Plesser, M.R., Roan, S.S., New construction of mirror manifolds: probing moduli space far from Fermat points, {\em Essays on Mirror manifolds}, editor S-T Yau, International Press 1992, 408--450.

%\bibitem{ghj}Gross, M., Huybrechts, D., Joyce, D., {\em Calabi-Yau manifolds and related geometries}, Springer, New York, 2003.

\bibitem{ph} Horja, P., Hypergeometric functions and mirror symmetry in toric varieties, e-print math.AG 9912109v3.

\bibitem{Karp}Karp, L. R., Quantum symmetries and exceptional collections, e-print arXiv:0811.3375v1.

\bibitem{Katz}Katz, N., Another look at the Dwork family, preprint.

\bibitem{mey} Meyer C., {\em Modular Calabi-Yau Threefolds}, Fields
Institute Monographs {\bf 22} AMS, Providence, RI, 2005.

\bibitem{mor}Morrison, D., R., Picard-Fuchs equations and mirror maps for hypersurfaces, {\em Essays on mirror manifolds} 241--264, Int. Press, Hong Kong, 1992.

\bibitem{mu} Mukai, S. {\em An introduction to invariants and moduli}, Cambridge studies in advanced mathematics {\bf 81}, Cambridge University Press, Cambridge, 2003.

\bibitem{RV} Rodriguez Villegas, F., Hypergeometric families of
Calabi-Yau manifolds, in {\em Proceedings of the workshop on
Calabi-Yau varieties  and mirror-symmetry}, Fields Institute,
Toronto, July 23-29, 2001, eds. Yui, N. Lewis, J.D., Fields Inst.
Commm. Series {\em 38} (2003), AMS, pp. 223--232.


\bibitem{shio}Shioda T., An explicit algorithm
for computing the Picard number of certain algebraic surfaces, {\em
Amer. J. Math.} {\bf 108} (1986) 415--432.

\bibitem{sch} Schimmrigk, R., http://www.th.physik.uni-bonn.de/Supplements/cy.html

%\bibitem{STW}Strominger, A., Witten, E., New manifolds for superstring compactification, {\em Comm. Math. Phys.} {\bf 101} (1985), no. 3, 341--361.


\bibitem{Yau}Yau, S.-T., Compact Three-dimensional Kaehler manifolds with zero Ricci curvature, in Proceedings of the Argonne Symposium on Anomalies, Geometry, Topology, eds. W. A. Bardeen and A. R. White (World Scientific, Singapore, 1985), pp. 495--406.

\end{thebibliography}

\documentclass[10pt]{amsart}

\usepackage{epsfig,amsmath,amsthm, amssymb}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{proposition}[theorem]{Proposition}

\newtheorem{corollary}[theorem]{Corollary}

\numberwithin{equation}{section}

\title{Quotients of hypersurfaces in weighted projective space}
\author{GILBERTO BINI}

\email{gilberto.bini@unimi.it} \curraddr{{\sc Dipartimento di
Matematica, Universit\`a degli Studi di Milano  \\ Via C. Saldini 50 -
20133 Milano  Italy}}

\thanks{{\em Mathematics Subject Classification 2000}: 14J32, 14J70}

\begin{document}

\maketitle

\begin{abstract} In \cite{bvg} some quotients of one-parameter families of Calabi-Yau varieties are related to the family of Mirror
Quintics by using a construction due to Shioda. In this paper, we
generalize this construction to a wider class of varieties. More
specifically,  let $A$ be an invertible matrix with non-negative
integer entries. We introduce varieties $X_A$ and $\overline{M}_A$
in weighted projective space and in ${\mathbb P}^n$, respectively.
The variety $\overline{M}_A$ turns out to be a quotient of a Fermat
variety by a finite group. As a by-product, $X_A$ is a quotient of a
Fermat variety and $\overline{M}_A$ is a quotient of $X_A$ by a
finite group. We apply this construction to some families of
Calabi-Yau manifolds in order to show their birationality.
\end{abstract}

\section{Introduction}\label{sec1}

Hypersurfaces in weighted projective space have been investigated by
many authors in connection with Physics, in particular Mirror
Symmetry: see, for instance, \cite{cdk}, \cite{cls} and
\cite{Yau}. The usual quintic threefold is an example of
Calabi-Yau manifold in ordinary projective space. Its mirror can
be described in terms of quotients of a one-parameter family of
quintics, the Dwork pencil. In \cite{bvg}, we investigated other
one-parameter families of Calabi-Yau manifolds and related them to
the family of Mirror quintics. Our main tool was a construction due
to Shioda \cite{shio}, which clarified previous work in \cite{dgj}.

In the present paper we generalize this construction to
hypersurfaces in weighted projective space. Let $A$ be an invertible
matrix of size $n$ with non-negative integer entries. For such a
matrix we define a weighted homogeneous polynomial $F_A$, a sum of
$n$ monomials. Its zero locus gives a projective scheme $X_A$ in
weighted projective space $W{\mathbb P}^{n-1}(q_1, \ldots, q_n)$.
The weights are determined by the relation $A{\bf q}=d {\bf e}$,
where ${\bf q}=(q_1, \ldots, q_n)$, ${\bf e }=(1,1, \ldots, 1)$ and
$d$ is the smallest positive integer such that $dA^{-1}$ has integer
entries. If the total degree of $F_A$ equals the degree of the
anticanonical bundle and the singularities of $X_A$ are canonical,
then $X_A$ is a Calabi-Yau manifold as explained in Section
\ref{sec2}. It is therefore interesting to study the quotients of
$X_A$ in relation with the mirror varieties of $X_A$. For this
purpose, we introduce a manifold $\overline{M}_A \subset {\mathbb
P}^n$, which we refer to as the Shioda quotient. It is the image of
a suitable map (first introduced by Shioda in \cite{shio} and
applied to a similar context in \cite{bvg}) of a Fermat variety.
Actually, we prove that the Shioda quotient is the quotient of a
Fermat variety by a finite group. As a by-product, it is also the
quotient of $X_A$ by a finite group. We describe these groups in
detail.

It is natural to investigate deformation families of $X_A$ defined
by $F_{A,t}=F_A-tx_1\ldots x_n$, where  $x_i$ are variables of
degree $q_i$. They are particularly meaningful when the Hodge number
$h^{2,1}$ is one, since then $F_{A,t}$ gives a versal deformation. Some examples of deformation families of $X_A$
when $X_A$ is a Calabi-Yau manifold in weighted projective space are
given in \cite{sch}. Their mirror all have $h^{2,1}=1$. We investigate the birational classes associated to these families and prove they are birational to the mirror families of the first
four hypergeometric Calabi-Yau families studied by Rodriguez Villegas \cite{RV}
and listed in \cite{mey}, p. 134.

\section{Preliminaries}\label{sec2}

\subsection{Weighted projective spaces} We briefly recall the definition of weighted projective space. Let $(q_1, \ldots, q_n)$ be a sequence of positive integers. As customary, set
$$
W{\mathbb P}^{n-1}(q_1, \ldots, q_n):= ({\mathbb C}^{n} \setminus \{0\})/\sim,
$$
where the equivalence relation is
$$
(x_1, \ldots, x_n) \sim (\lambda^{q_1}x_1, \ldots, \lambda^{q_n}x_n)
$$
whenever $\lambda \in {\mathbb C} \setminus \{0\}$. Recall that
$$
W{\mathbb P}^{n-1}(q_1, \ldots, q_n) \cong W{\mathbb
P}^{n-1}(q_1/a_1, \ldots, q_n/a_n),
$$
where $a_i=l.c.m.(d_1, \ldots, \widehat{d_i}, \ldots, d_n)$ and
$d_i=g.c.d.(q_1, \ldots, \widehat{q_i}, \ldots, q_n)$. Weighted
projective space are almost always singular. As proved in \cite{Di},
the singular locus of $W{\mathbb P}^{n-1}(q_1, \ldots, q_n)$ can be
described in the following way. Let $p$ be a prime. Let $I(x)=\{j;x_j \neq 0\}$. Then define
$$
Sing_p(W{\mathbb P}^{n-1}):=\{ x \in W{\mathbb P}^{n-1}(q_1,\ldots,
q_n) : p|q_i \, \hbox{for any}\, i \in I(x)\}.
$$

The singular locus of the weighted projective space is given by the
union over all primes of $Sing_p(W{\mathbb P}^{n-1})$.

 As explained, for instance, in \cite{CK},  weighted projective space is a toric variety.  Set $Q:=\sum_i q_i$. We recall that the canonical sheaf of $W{\mathbb P}^{n-1}(q_1, \ldots, q_n)$ is given by ${\mathcal O}( - Q)$, which is not always a line bundle. As proved in \cite{CK}, Lemma 3.5.6., the canonical sheaf is a line bundle if and only if $q_i |Q$ for all $i=1, \ldots, n$. Under this assumption, weighted projective space is a Fano toric variety.

Some of our hypersurfaces are Calabi-Yau varieties. Following \cite{CK}, a (possibly singular) Calabi-Yau variety is an $m$-dimensional normal compact variety $X$ which satisfies the following conditions: \\
(i) $X$ has at most Gorenstein canonical singularities; \\
(ii) the dualizing sheaf of $X$ is trivial;\\
(iii) $H^1(X, {\mathcal O}_X)=\ldots = H^{m-1}(X, {\mathcal O}_X)=0$.

  Let $f$ be a weighted homogeneous polynomial such that the zero locus $\{f=0\}$ is quasi-smooth (according to Definition 3.1.5 in \cite{Dol}).
  Since a quasi-smooth scheme has finite quotient singularities (\cite{Dol}, Thm. 3.1.6), the locus
   $\{f=0\}$ is normal and Cohen-Macaulay; furthermore, it is Gorenstein with dualizing sheaf ${\mathcal O}(d)$,
   where $d$ is the weighted degree of $f$. Assume that $\{f=0\}$ has at most canonical singularities. Following the arguments in
   Proposition 4.1.3 in \cite{CK}, it follows that $\{f=0\}$ is a Calabi-Yau variety when $d=Q$. Notice that if there exists a crepant
   resolution of $\{f=0\}$, the singularities are canonical by
   definition.

\subsection{The Shioda maps} Let $A$ be an invertible matrix with non-negative integer entries:
$$
A=(a_{ij}) \quad (\in M_{n}({\mathbb Z})), \quad a_{ij} \in {\mathbb Z}_{\geq 0}, \quad det(A) \neq 0.
$$

For such a matrix we define a polynomial in $n$ variables, a sum of $n$ monomials:
$$
F_A:=\sum_{i=1}^n\prod_{j=1}^n x_j^{a_{ij}}= x_1^{a_{11}}x_2^{a_{12}}\ldots x_n^{a_{1n}} + x_1^{a_{21}}x_2^{a_{22}}\ldots x_n^{a_{2n}} + \ldots
$$

Let $d$ be the smallest positive integer such that $B:=dA^{-1}$ is in $M_n({\mathbb Z})$. Then set
\begin{equation}
\label{qu}
{\bf q}:= B{\bf e},
\end{equation}
where ${\bf e}=(1, \ldots, 1)$ and ${\bf q}=(q_1, \ldots, q_n)$. Clearly, this implies that
\begin{equation}
\label{di}
A{\bf q}=d{\bf e}.
\end{equation}

The zero locus $X_A=Z(F_A)$  is a (not necessarily smooth or irreducible) projective variety $X_A$, which is contained
in $W{\mathbb P}^{n-1}(q_1, \ldots, q_n)$. Let $m$ be the greatest common divisor of the $q_i$'s. Define $a_i=q_i/m$.
Thus we have $$X_A \subset W{\mathbb P}^{n-1}(q_1, \ldots, q_n) \cong W{\mathbb P}^{n-1}(a_1, \ldots, a_n).$$

We have a rational map $\phi_A$ from ${\mathbb P}^{n-1}$ of degree $d$ to $X_A$ defined by:
$$
\phi_A: {\mathbb P}^{n-1} \rightarrow W{\mathbb P}^{n-1}(q_1, \ldots, q_n),$$
  $$(y_1: \ldots : y_n) \rightarrow (x_1: \ldots : x_n), \quad x_j=\prod_{k=1}^ny_k^{b_{jk}}.
 $$

Notice that each $y_j$ has degree one, so $deg(x_j) = \sum_k b_{jk}=q_j$. Hence $\phi_A$ is indeed a rational map from ${\mathbb P}^{n-1}$
to $W{\mathbb P}^{n-1}(q_1, \ldots, q_n)$.

Assume further that $d=Q=\sum_jq_j$, so that the variety $X_A$ gives
a Calabi-Yau provided the singularities are canonical.  We can read
this condition on the coefficients of the matrix $A^{-1}$. In fact,
we have:
$$
d= Q= \sum_j q_j ={}^t{\bf e}{\bf q}= d ^t{\bf e}A^{-1}{\bf e},
$$
which gives
\begin{equation}
\label{boh}
^t{\bf e}A^{-1}{\bf e}=1;
\end{equation} in other words, $\sum_{ij}a_{ij}'=1$, where $a'_{ij}$ are the entries of $A^{-1}$.

We define a rational map
$$
q_A:W{\mathbb P}^{n-1}(q_1, \ldots, q_n) \rightarrow \overline{Im(q_A)}:=\overline{M}_A \subset {\mathbb P}^n
$$
in the following way:
$$
(x_1:\ldots :x_n) \rightarrow (u_0: u_1 : \ldots : u_n):=\left(\prod _{j=1}^n x_j : \prod_{j=1}^nx_j^{a_{1j}}:
 \ldots : \prod_{j=1}^n x_j^{a_{nj}}\right).
$$

Since $deg(x_i)=q_j$ and $A{\bf q}= d{\bf e}$, we have
$$
deg(u_0)=\sum_j q_j=d, \qquad deg(u_k)=\sum_j a_{kj} q_j =d,
$$
hence $q_A$ is well-defined.

Finally, we describe the composition $q_A \circ \phi_A$, which will be used in the next section. First, as $BA=AB=dI_n$ (where $I_n$ is the identity matrix) we have $\sum_j a_{lj}b_{jk}=d\delta_{lk}$, where $\delta_{kl}$ is the Kronecker delta.
 Second, we set
$$
{\bf q}':=d\,{} ^tA^{-1}{\bf e}= d \, ^t {\bf e} A^{-1} = \, ^t {\bf e} B;
$$
so $q_k'=\sum_j b_{jk}.$ This said, it is easy to check that the composition $q_A \circ \phi_A: X_d \subset {\mathbb P}^{n-1} \rightarrow {\overline M}_A \subset {\mathbb P}^n$ is given by
\begin{equation}
\label{composition} (u_0: u_1 : \ldots : u_n) = \left(
\prod_{k=1}^ny_k^{q'_k}: y_1^d : \ldots : y_n^d \right),
\end{equation}
where $X_d$ is the Fermat variety $\{ \sum_{i=1}^n y_i^d=0\}$ and
${\overline M}_A$ is the closure of the image of $X_d$.

Let us consider the projection
\begin{equation}
\label{forget}
\begin{array}{lcr}
\pi: {\overline M}_A \subset {\mathbb P}^n & \rightarrow & V \cong {\mathbb P}^n \\
(u_0: u_1: \ldots : u_n) & \rightarrow & (u_1 : \ldots : u_n),
\end{array}
\end{equation}
where $V$ is the closure of $\pi({\overline M}_A)$.  It is easy to check
that $V$ is isomorphic to the ${\mathbb P}^{n-1}$ given by
$\sum_{i=1}^n u_i=0$ as the Fermat equation has to be satisfied.

From now on, we will assume that $q_k'$  is strictly positive for
any $k$. By direct inspection, we obtain the following equations for
the image of the Fermat variety $X_d$ under $q_A \circ \phi_A$:
\begin{equation}
u_1 + \ldots + u_n =0, \quad u_0^d= u_1^{q_1'} \ldots u_n^{q_n'}.
\label{first}
\end{equation}

Set $m':=g.c.d(d, q_1', \ldots, q_n')$. Hence $d=m'a'$ and
$q_k'=a_k'm'$, so the composition $q_A \circ \phi_A$ is given by
\begin{equation}
\label{newim}
(u_0:\ldots:u_n)=\left(\prod_{k=1}^ny_k^{q'_k}: y_1^d : \ldots : y_n^d \right)=\left(
\prod_{k=1}^n(y_k^{m'})^{a'_k}: (y_1^{m'})^{a'} : \ldots : (y_n^{m'})^{a'} \right).
\end{equation}

By composing \eqref{newim} with the map $t_k=y^{m'}_k$ for $k=1, \ldots, n$, we get
$$
(u_0:\ldots:u_n)=\left(\prod_{k=1}^nt_k^{a'_k}: t_1^{a'} : \ldots : t_n^{a'} \right),
$$
so the equations defining ${\overline M}_A$ are the following:
\begin{equation}
\label{second}
u_1+ \ldots + u_n=0, \quad u_0^{a'}= u_1^{a_1'} \ldots u_n^{a_n'}.
\end{equation}

 In the next section, we will show that under our assumptions on the $q_j$'s, the equation \eqref{second} define a very singular  model in $(n-1)$-dimensional projective space for  a manifold with $h^{n-2,0}=1$ of degree $a' ( > n$ in general).

\section{The Shioda quotient}\label{sec3}

In this section we assume that $A \in M_n({\mathbb Z}_{\geq 0})$ is an invertible matrix such that \eqref{boh} holds. We will introduce "natural" automorphism groups and study the quotients by these groups.

\subsection{The automorphism groups}
 Let $\zeta= \zeta_d$ be a generator of the cyclic group of $d$-th roots of unity, where $d$ is the smallest positive integer such that $dA^{-1}$ has integer entries.
For ${\bf k}=(k_1, \ldots, k_n) \in ({\mathbb Z}/d{\mathbb Z})^n$ we define an automorphism $g_{{\bf k}}$ of ${\mathbb P}^{n-1}$ by
$$
g_{{\bf k}}(y_1:\ldots : y_n):=\left(\zeta^{k_1}y_1 : \ldots : \zeta^{k_n} y_n\right).
$$
Note that ${\bf a},{\bf b} \in \mu_d^n$ define the same automorphism iff ${\bf a} - {\bf b} =
(k, \ldots, k)$ for some $k \in \mu_d$. Define $\Gamma_d$ to be the
quotient group
$$
\Gamma_d:=\mu_d^n/\langle g_{(1,1, \ldots, 1)} \rangle  \, \, \left ( \subset Aut({\mathbb P}^{n-1})\right)
$$

Notice that $\Gamma_d \cong \mu_d^{n-1}$, hence $\# \Gamma_d = d^{n-1}$. The group $\Gamma_d$ is a
subgroup of the automorphism group of the Fermat variety $X_d$.

The map $q_A \circ \phi_A$ \eqref{composition} is invariant under
the subgroup of $\Gamma_d$ given by
$$
\Gamma({\bf q'}):=\left\{ g_{{\bf k}}: {\bf k}= (k_1, \ldots,
k_n); \sum_j k_j q'_j \equiv 0 \, \, \hbox{mod} \, \, {d}\right\}\Biggl/ \langle g_{(1,\ldots, 1)} \rangle.
$$

In other words, we have
$$
(q_A \circ \phi_A)(g_{{\bf k}}(y_1: \ldots: y_n))=(q_A \circ \phi_A)
(y_1: \ldots: y_n)
$$
for all $g_{{\bf k}}$ in $\Gamma({\bf q}')$ and all $(y_1, \ldots, y_n) \in X_d$.
Notice that $g_{(1, \ldots, 1)}$ is an element of $\Gamma({\bf q'})$. In fact, by \eqref{boh}:
$$
\sum_j q'_j= \, ^t{\bf e}{\bf q}'=d\, ^t{\bf e}\, ^tA^{-1}{\bf e}=d.
$$

The coordinate functions of the Shioda map
$\phi_A$ are products of the $y_i$. If $\phi_A(y_1: \ldots: y_n)=(x_1: \ldots : x_n)$, then
$$
\phi_A(g_{{\bf k}}(y_1: \ldots: y_n))= \left( \zeta^{k'_1}x_1:
\ldots : \zeta^{k'_n}x_n \right).
$$

As $x_j= \prod y_k^{b_{jk}}$, the column vector ${\bf k}'$ is
obtained from the column vector ${\bf k}$ as ${\bf k}'=B{\bf
k}$. Thus we get a homomorphism
$$
\Gamma({\bf q}') \rightarrow Aut(X_A), \quad g_{\bf k} \rightarrow g_{B{\bf k}},
$$
which is well defined since $B{\bf e} \equiv 0$ mod $d$.

The kernel (image resp.) of this homomorphism will be denoted by $\Gamma_A$ ($H_A$ resp.).
Notice that $\Gamma_A$ is the subgroup of $\Gamma({\bf q}')$, which is generated by the images of the $g_{{\bf k}}$
such that $B{\bf k} \equiv 0$ mod $d$.

\subsection{The birational model}   We recall that two rational maps between
algebraic varieties $f_i: X \rightarrow Y_i$ for $i=1,2$ are said to be birationally
 equivalent if there is a Zariski open subset $U$ of $X$ and
 there are Zariski open subsets $U_i \subset Y_i$ with
 an isomorphism $\phi: U_1 \rightarrow U_2$ such that $\phi \circ f_1 = f_2$ on $U$.

\begin{theorem}
\label{central}
Let $A$ be an invertible $n \times n$ matrix with integer entries such that $X_A$ is irreducible and \eqref{boh} holds. Then the
composition $q_A \circ \phi_A$ is birational to the quotient map $X_d \rightarrow X_d/\Gamma({\bf q}')$; hence $X_d/\Gamma({\bf q'})$ is birational to ${\overline M}_A$.
\end{theorem}
\proof

The composition $q_A \circ \phi_A$ is given by \eqref{composition}.
Also, recall the map $\pi$ defined in  \eqref{forget}. The
composition of $ q_A \circ \phi_A $ and $\pi$ yields a map from the
Fermat variety $X_d$ to $V \cong {\mathbb P}^{n-1}$, which
corresponds to an abelian extension  with group $\Gamma_d$ of
function fields - recall that $X_d$ is the Fermat variety and $u_i =
y_i^d$. This means that $X_d \rightarrow V$ is the quotient for the
group $\Gamma_d$, namely $X_d/\Gamma_d = V$. Therefore, by abelian
Galois theory, each subfield is obtained as an invariant field under
a finite subgroup of $\Gamma_d$. Thus, the map $X_d \rightarrow
\overline{M}_A$ corresponds to a quotient by a finite subgroup of
$\Gamma_d$. Now, we show that this subgroup is isomorphic to
$\Gamma({\bf q}')$. The map $\pi$ corresponds to an abelian
extension of function fields with group ${\mathbb Z}/a'{\mathbb Z}$,
where $d=m'a'$ and $m'=g.c.d.(d, q_1', \ldots, q_n')$.  The group
$\Gamma_d$ acts on ${\overline M}_A$ only through the variable
$u_0$, and $g_{\bf k}:u_0 \rightarrow \zeta_{m'}^{{\bf q}'\cdot {\bf
k}}u_0$, where $\zeta_{m'}=\zeta^{a'}$ is a primitive $m'$-th root
of unity. The kernel of this action is exactly $\Gamma({\bf q'})$,
hence the map $X_d \rightarrow {\overline M}_A$ corresponds to an
extension of function fields with group $\Gamma({\bf q'})$. Hence,
the claim follows.
\endproof

By the Theorem above, the order of the group $\Gamma({\bf q}')$ is $d^{n-2}m'$.

\begin{corollary}
The maps $\phi_A : X_d \rightarrow X_A$ and $q_A: X_A \rightarrow \overline{M}_A$ are birational to quotient maps. In particular, $X_A$ (${\overline M}_A$ resp.) is birational to $X_d/\Gamma_A$ ($X_A/H_A$ resp.).
\end{corollary}

\proof By Theorem \ref{central}, the composition of $\phi_A$ and $q_A$ is a quotient map, namely:
$$
X_d \rightarrow X_A \rightarrow {\overline M}_A \approx X_d/\Gamma({\bf q}').
$$

The proof follows easily from arguments similar to those in Theorem 2.6 in \cite{bvg}.
\endproof

Now, we prove that the equations \eqref{first} and \eqref{second} give a very singular model for a manifold with $h^{n-2,0}=1$.
Assume, $q_i' > 0$. We recall that the vector space of holomorphic $(n-2)$-forms on a smooth
hypersurface $X=Z(F)$ of degree $d \geq n$ in ${\mathbb P}^{n-1}$ has a basis of the form
$$
\omega_{{\bf b}, F}=Res_X \left(y_1^{b_1}\ldots y_n^{b_n} \frac{\sum_i ^n (-1)^i y_i dy_1 \wedge \ldots \wedge \widehat{dy_i}
\wedge \ldots \wedge dy_n}{F} \right),
$$
where ${\bf b}=(b_1, \ldots, b_n)$ and $b_i \in {\mathbb Z}_{\geq 0}$, $\sum_i b_i = d-n$.

\begin{proposition}
There exists a unique holomorphic $(n-2)$ form on any resolution of ${\overline M}_A$.
\end{proposition}

\proof The action of an element $g_{{\bf k}}$ in $\Gamma({\bf q'})$
on $\omega_{{\bf b}, F}$ is given by $\zeta^{\sum_i (b_i+1)k_i}$,
where $b_i \in {\mathbb Z}_{\geq 0}$. A form is invariant with
respect to $\Gamma({\bf q'})$ if and only if $b_i+1=q'_i$ for all
$i$. The unique invariant form $\omega_{{\bf b}, F}$ with vector
${\bf b}={\bf q'}- {\bf e}$ descends to a form on the quotient
$X_d/\Gamma({\bf q'}) \approx {\overline M}_A$.
\endproof

\subsection{Some examples}\label{sec7}

{\it Example A.}  Let us consider
$F_A:=x_1^5+x_2^{10}+x_3^{10}+x_4^{10}+x_5^2=0$  in weighted
projective space $W{\mathbb P}^4(2,1,1,1,5)$. It is easy to check
that the corresponding hypersurface is smooth and does not intersect
the singularities of $W{\mathbb P}^4(2,1,1,1,5)$, which are two
isolated points.

The  matrix $A$ is given by $diag(5,10,10,10,2)$. The matrix $B$ is $diag(2,1,1,1,5)$ and $d=10$. The condition ${}^t{\bf
e}A^{-1}{\bf e}=1$ is satisfied. Moreover,
$$
{\bf q}=(2,1,1,1,5), \quad {\bf q}'=(2,1,1,1,5)
$$
so $q_i' >0$ for any $i=1, \ldots, 5.$ The equations cutting out
$\overline{M}_A$ in ${\mathbb P}^5$ are given by
$$
u_1+u_2+u_3+u_4+u_5=0,
$$
$$
u_0^{10}=u_1^2u_2u_3u_4u_5^5.
$$

The integer $d = 10$. Generators for the groups $\Gamma({\bf q}')$, $\Gamma_A$ and $H_A$ are as follows. Consider the elements ${\bf v_1}, \ldots , {\bf v_4} \in ({\mathbb Z}/10{\mathbb Z})^5$ given by:
$$
{\bf v}_1=(0,0,0,5,1), \qquad {\bf v}_2=(0,0,1,4,1),
$$
$$
{\bf v}_3=(1,0,0,3,1), \qquad {\bf v}_4=(0,1,0,4,1),
$$
then $^t {\bf q'}\cdot {\bf v}_i \equiv 0$ mod $10$ for $i=1, \ldots, ,4$. Moreover, we have ${\bf e}= 8{\bf v}_1+{\bf v}_2+{\bf v}_3+{\bf v}_4$; hence
$$
\Gamma({\bf q}') = \langle {\bf v}_1, {\bf v}_2, {\bf v}_3 \rangle \cong \mu_{10}^3.
$$

The group $\Gamma_A$ is isomorphic to $\mu_{10}$ and is generated by $(5,0,0,0,6)$. Finally, set
$$
{\bf w}_1=(0,0,1,4,5), \quad {\bf w}_2=(2,0,0,3,5), \quad {\bf w}_3=(0,1,0,4,5).
$$

As ${\bf w}_1 + {\bf w }_2 + {\bf w}_3 = {\bf q}$, we have
$$
H_A= \langle {\bf w}_1, {\bf w}_2 \rangle \cong \mu_{10}^2 \subset
Aut(X_A).
$$

The isomorphism $H_A \cong \mu_{10}^2$ was first suggested in
\cite{gpr}.  Finally, as mentioned in \cite{mor}, notice that
$h^{2,1}(X_A)=1$.

{\it Example B.} Let us consider the equation
$F_A:=x_1^{15}x_5+x_2^5+x_3^5+x_3x_4^5+x_2x_4^2=0$ in weighted
projective space $W{\mathbb P}^4(1,5,5,4,10)$. In this case, we do
not have a Fano toric veariety since $4$ does not divide
$25=1+5+5+4+10$. The zero locus does not intersect the singularities
of $W{\mathbb P}^4(1,5,5,4,10)$ and $F_A$ is a smooth variety. The
matrices $A$ and $B$ are given by
$$
A:=\left(
\begin{array}{ccccc}
15 & 0 & 0 & 0 & 1 \\
0 & 5 & 0 & 0 & 0 \\
0 & 0 & 5 & 0 & 0 \\
0 & 0 & 1 & 5 & 0 \\
0 & 1 & 0 & 0 & 2
\end{array}
\right),
 \quad B:= \left(
\begin{array}{ccccc}
10 & 1 & 0 & 0 & -5 \\
0 & 30 & 0 & 0 & 0 \\
0 & 0 & 30 & 0 & 0 \\
0 & 0 & -6 & 30 & 0 \\
0 & -15 & 0 & 0 & 75
\end{array}
\right).
$$

The integer $d$ equals $150$. Moreover, we have
$$
{\bf q}=6(1, 5, 5, 4, 10), \qquad {\bf q}'=(10,16,24,30,70)
$$

and $a'=g.c.d.(d, q_1', \ldots, q_5')=2$. Notice that $q_i' >0$ for any $i=1, \ldots, 5$.

A birational model of ${\overline M}_A$ is cut out by the equations
\begin{equation}
\label{notrans}
u_0^{75}=u_1^5u_2^8u_3^{12}u_4^{15}u_5^{35}, \quad u_1 + \ldots + u_5=0.
\end{equation}

Consider the vectors ${\bf r_i} \in {\mathbb Z}/d{\mathbb Z}$ given by:
$$
{\bf r}_1= (0,0,75,0,0), \qquad {\bf r}_2=(0,1,1,0,8),
$$
$$
{\bf r}_3=(1,0,0,0,2), \quad {\bf r}_4=(0,0,0,1,6), \quad {\bf r}_5=(0,0,5,0,9);
$$
then $9{\bf e}={\bf r}_1+9 {\bf r}_2+ 9{\bf r}_3 + 9{\bf r}_4- 15 {\bf r}_5$. The vectors ${\bf r}_i$ generate $\Gamma({\bf q}')$; in fact, the following holds:
$$
\Gamma({\bf q}')\cong \langle {\bf r}_1, {\bf r}_2, {\bf r}_3, {\bf r}_4, {\bf r}_5 \rangle/ < {\bf r}_1+9 {\bf r}_2+ 9{\bf r}_3 + 9{\bf r}_4- 15 {\bf r}_5 >\cong {\mathbb Z}/2{\mathbb Z} \times \left( {\mathbb Z}/150{\mathbb Z}\right)^3
$$

By using Magma, it is possible to check that
$$
\Gamma_A \cong \left( {\mathbb Z}/150{\mathbb Z}\right)^3,
$$
$$
H_A \cong {\mathbb Z}/2{\mathbb Z}.
$$

The group $H_A$ is generated by $g_{(0,75,75,0,75)}$, which maps
$(x_1: \ldots : x_5)$ to $(x_1: - x_2: - x_3 : x_4: -x_5)$. Recall
that $X_A \rightarrow {\overline M}_A \approx X_A/H_A$ is a double
cover.

 {\it Example C}. Consider the weighted homogeneous polynomial  $${F_A}:=x_1^2+x_2^3+x_3^{18}+x_4^{18}+x_5^{18}$$. It gives a quasi-smooth locus in weighted projective space $W{\mathbb P}^4(9,6,1,1,1)$.
 The matrices $A$ and $B$ are given by $diag(2,3,18,18,18)$ and
 $diag(9,6,1,1,1)$, respectively.

The integer $d$ is equal to $18$ and we have
$$
{\bf q}= {\bf q}'=(9,6,1,1,1).
$$

Notice that $q_i' >0$ for any $i$. Moreover, using Magma we found that
$$
\Gamma({\bf q}) \cong \left( {\mathbb Z}/18{\mathbb Z}\right)^3;
$$
$$
H_A \cong {\mathbb Z}/6{\mathbb Z} \times {\mathbb Z}/18{\mathbb Z};
$$
$$
\Gamma_A \cong {\mathbb Z}/3{\mathbb Z} \times {\mathbb Z}/18{\mathbb Z}.
$$

The Calabi-Yau $X_A$ has a singularity at $[-1,1,0,0,0]$, which is a singularity of $W{\mathbb P}^4(9,6,1,1,1)$. As explained in \cite{Karp}, this singularity can be blown-up so as to get a Calabi-Yau in the (toric) blow-up of $W{\mathbb P}^4(9,6,1,1,1)$.

{\it Example D}. When the group $H_A$ is trivial, it is possible to
write down an explicit birational inverse between from
$\overline{M}_A$ to $X_A$. We show it in one specific example. Let us
consider the polynomial
$$
F_A:=x_1^5+x_2^9x_3+x_3^9x_4+x_4^{10}+x_5^{2}
$$
where $A$ is the matrix $$
A:=\left(
\begin{array}{ccccc}
5 & 0 & 0 & 0 & 0 \\
0 & 9 & 1 & 0 & 0 \\
0 & 0 & 9 & 1 & 0 \\
0 & 0 & 0 & 10 & 0 \\
0 & 0 & 0 & 0 & 2
\end{array}
\right).
$$

The variety $X_A:=Z(F_A)$ is contained in $W{\mathbb P}^4(2,1,1,1,5)$. The matrix $B$ is given by
$$
B:=\left(
\begin{array}{ccccc}
162 & 0 & 0 & 0 & 0 \\
0 & 90 & -10 & 1 & 0 \\
0 & 0 & 90 & -9 & 0 \\
0 & 0 & 0 & 81 & 0 \\
0 & 0 & 0 & 0 & 405
\end{array}
\right),
$$
where $AB=BA=810I$. The map $q_A: W{\mathbb P}^4(2,1,1,1,5)
\rightarrow {\mathbb P}^5$
$$
(x_1: \ldots : x_5) \rightarrow (u_0: \ldots : u_5):= (x_1x_2\ldots
x_5: x^5_1: x_2^9x_3: x_3^9x_4:x_4^{10}x_2:x_5^{2})
$$
maps $X_A$ to the variety
$$
\overline{M}_A:=Z(u_1+\ldots+u_5,
-u^{810}+u_1^{162}u_2^{90}u_3^{80}u_4^{73}u_5^{405}) \subset
{\mathbb P}^5.
$$

An explicit birational inverse for $q_A$ is given by the following
map:

$$
\left\{
\begin{array}{l}
M^{162}x_1=u_1^{65}u_2^{54}u_3^{12}u_4^{15}u_5^{162}, \\
\\
M^{81}x_2=u_1^{-2}u_2^8u_3^{-11}u_4^{-8}u_5^{-5}, \\
\\
M^{81}x_3=u_1^{18}u_2^{19}u_3^{-1}u_4u_5^{45}, \\
\\
M^{81}x_4=u_2^9u_3^{-10}u_4^{-7}, \\
\\
M^{405}x_5=u_1^{81}u_2^{90}u_3^{-10}u_4u_5^{203},
\end{array}
\right.
$$
where $M=x_1^2x_2^3x_3x_4x_5^2$.

\section{A one-dimensional family}\label{sec4}

Let us consider the one-parameter family ${\mathcal X} \rightarrow {\mathbb P}^1_t$ of degree $d$ hypersurfaces
in ${\mathbb P}^{n-1}$ with ${\mathcal X}_t=X_{d,t}=Z(F_{d,t})$, where
\begin{equation}
\label{mirror}
F_{d,t}= \sum_i^n y_i^d - t \, y_1^{q_1'}\ldots y_n^{q'_n}.
\end{equation}

Clearly, this is a one-dimensional deformation of the Fermat variety $X_d$.  If we apply the map $\phi_A: X_{d,t} \subset
{\mathbb P}^{n-1} \rightarrow X_{A,t} \subset W {\mathbb P}(q_1, \ldots, q_n)$, the image of $X_{d,t}$ is given by
$
X_{A,t}=Z(F_{A,t}),
$ where
$$
F_{A,t} = F_A- t\, x_1x_2\ldots x_n.
$$

Under the composition $q_A \circ \phi_A$ the image
 of \eqref{mirror} is given by the equations
 \begin{equation}
 \label{uzero}
\sum_i u_i - t u_0=0, \qquad u^d_0 = u_1^{q_1'}\ldots u_n^{q_n'}.
\end{equation}

If $t \neq 0$, we solve for $u_0$ and get the equation
\begin{equation}
\label{mirr}
\left( \sum_i ^n u_i \right)^d = t^d u_1^{q_1'} \ldots u_n^{q_n'}.
\end{equation}

The group $\Gamma({\bf q'})$ acts on each $X_{A,t}$ since $\sum_i q'_i$ equals $d$. Denote by $\overline{\mathcal M} \rightarrow {\mathbb P}^1_t$ the family
given by \eqref{mirr}. By the universal property of the quotient there exists a map $\Psi$ between
${\mathcal X}/\Gamma({\bf q'})$
and ${\overline {\mathcal M}}$, which commutes with the projection map on ${\mathbb P}^1_t$.

\begin{proposition}
The map $\Psi$ yields a birational morphism from ${\mathcal X}/ \Gamma({\bf q}')$ to ${\overline{\mathcal M}}_t$.
\end{proposition}
\proof It suffices to compare the degree of the quotient map ${\mathcal X} \rightarrow {\mathcal X}/\Gamma({\bf q'})$, which is $\#\Gamma({\bf q'})$, with that
of the map ${\mathcal X} \rightarrow \overline{{\mathcal M}}$. Let
$$
(l_0:l_1 \ldots: l_n)=(\prod_k^ny_k^{q_k'}: y_1^d : \ldots y_n^d)
$$
be a generic point in $\overline{Im( q_A \circ \phi_A)}$. Thus, we have $y_j=\zeta^{k_j}\sqrt[d]{l_j}$, where $\zeta$ is a primitive
 $d$-th root of unity. Hence, we get
 $$
 \zeta^{\sum_j q_j'k_j}\sqrt[d]{l_1^{q_1'}\ldots l_n^{q_n'}}=l_0.
 $$

On the other hand, by the equation of $F_{A,t}$,  we must have
\begin{equation}
\label{set}
 \sum_j q_j'k_j \equiv 0 \, \, \hbox{mod} \, \,  d.
\end{equation}

Recall that $\sum_j q_j' \equiv 0$ mod $d$, so we can take the quotient of the set \eqref{set} of solutions
by $(1, 1, \ldots, 1)$. The degree of ${\mathcal X} \rightarrow {\overline {\mathcal M}}$ is thus equal to $\# \Gamma({\bf q'})$.

\endproof

\subsection{Some Birational Families}\label{sec8} Let us examine some Calabi-Yau varieties $X_A$ which have a one-dimensional versal deformation family,
so it may be described by the family in the section above.  In
Schimmrigk's list (see \cite{sch}), we found twelve entries with
$h^{2,1}=1$.  Since the deformation space is one dimensional, we
take into account one-dimensional families corresponding to these
entries of the list. The generic members of the families, which are denoted by
the same letter, have the same Euler characteristic. We have

$$
{\mathcal
A}_1(t)=x_1^8x_3+x_2x_3^7+x_2^7x_4+x_4^7+x_5^2-tx_1x_2x_3x_4x_5
\subset {\mathbb A}^1 \times W{\mathbb P}^4(75,84,86,98,343),
$$
$$
{\mathcal
A}_2(t)=x_1^8x_2+x_2^7x_3+x_3^7+x_4^4+x_4x_5^2-tx_1x_2x_3x_4x_5
\subset {\mathbb A}^1 \times W{\mathbb P}^4(43,48,56,98,147),
$$
$$
{\mathcal
A}_3(t)=x_2^8+x_1^7x_3+x_3^7+x_1x_4^4+x_4x_5^2-tx_1x_2x_3x_4x_5,
\subset {\mathbb A}^1 \times W{\mathbb P}^4(48,49,56,86,153)$$
$$
{\mathcal
A}_4(t)=x_1x_2^7+x_1^7x_3+x_3^7+x_2x_4^4+x_4x_5^2-tx_1x_2x_3x_4x_5\subset
{\mathbb A}^1 \times W{\mathbb P}^4(42,43,49,75,134),
$$

$$
{\mathcal
B}_1(t)=x_1^{10}x_2+x_2^9x_3+x_3^9+x_4^5+x_5^2-tx_1x_2x_3x_4x_5
\subset {\mathbb A}^1 \times W{\mathbb P}^4(73,80,90,162,405),
$$
$$
{\mathcal
B}_2(t)=x^9_1x_2+x_2^9+x_1x_3^8+x_3x_4^5+x_5^2-tx_1x_2x_3x_4x_5
\subset {\mathbb A}^1 \times W{\mathbb P}^4(64,72,73,115,324),
$$
$$
{\mathcal
B}_3(t)=x_1^9x_2+x_2^9+x_1x_3^5+x_4^5+x_3x_5^2-tx_1x_2x_3x_4x_5\subset
{\mathbb A}^1 \times W{\mathbb P}^4(40,45,73,81,166),
$$
$$
{\mathcal
B}_4(t)=x_1^9+x_2^8+x_2x_3^5+x_1x_4^5+x_4x_5^2-tx_1x_2x_3x_4x_5\subset
{\mathbb A}^1 \times W{\mathbb P}^4(40,45,63,64,148),
$$

$$
{\mathcal
C}_1(t)=x_1^6x_3+x_2x_3^5x_2^5+x_4+x_4^5+x_5^3-tx_1x_2x_3x_4x_5\subset
{\mathbb A}^1 \times W{\mathbb P}^4(52,60,63,75,125),
$$
$$
{\mathcal
C}_2(t)=x_2^6+x_1^5x_3+x_3^5+x_1x_4^4+x_4x_5^3-tx_1x_2x_3x_4x_5
\subset {\mathbb A}^1 \times W{\mathbb P}^4(48,50,60,63,79),
$$

$$
{\mathcal
D}_1(t)=x_1^5x_3+x_3^4x_4+x_2x_4^4+x_2^4x_5+x_5^4-tx_1x_2x_3x_4x_5\subset
{\mathbb A}^1 \times W{\mathbb P}^4(41,48,51,52,64),
$$
$$
{\mathcal
D}_2(t)=x_3^5+x_1^5x_4+x_2x_4^4+x_2^4x_5+x_5^4-tx_1x_2x_3x_4x_5\subset
{\mathbb A}^1 \times W{\mathbb P}^4(51,60,64,65,80).
$$

Let $V_5(t), V_6(t), V_8(t), V_{10}(t)$ be the four hypergeometric families on page 134
in \cite{mey}. For each of these families $V_j(t)$, $j=5,6,8,10$,
there is a group acting on the family such that the mirror $W_j(t)$
of $V_j(t)$ can be described as a resolution of the quotient
$V_j(t)/G$. The singular members have one orbit of ordinary nodes
under the action of $G$ and the resolution of the quotient is a
rigid Calabi-Yau threefold, i.e., $h^{2,1}=0$.

\begin{theorem}
The following birational equivalences hold:
$$
{\mathcal A_1}(t)  \approx {\mathcal A_2}(t) \approx {\mathcal
A_3}(t)\approx {\mathcal A_4}(t) \approx W_8(t),
$$
$$
{\mathcal B_1}(t)  \approx {\mathcal B_2}(t) \approx {\mathcal
B_3}(t)\approx {\mathcal B_4}(t) \approx W_{10}(t),
$$
$$
{\mathcal C_1}(t)  \approx {\mathcal C_2}(t) \approx W_6(t),
$$
$$
{\mathcal D_1}(t)  \approx {\mathcal D_2}(t) \approx W_5(t).
$$
\end{theorem}

\proof It is easy to check that the general member of the families
above is a singular Calabi-Yau in four weighted projective space.
The singular locus is a rational curve. For some values of $t$ there
are extra singularities that are ordinary nodes, namely:
\begin{center}
\begin{tabular}{cccc}
${\mathcal A}_i(t)$ & $t^8=2^{16}$ \vline & ${\mathcal B}_i(t)$ &
$t^{10}=800000$ \\ \hline ${\mathcal C}_i(t)$ & $t^6=3^62^4$ \vline
& ${\mathcal D}_i(t)$ & $t^5=5^5$ \\ \hline
\end{tabular}
\end{center}

Let us focus on the two families ${\mathcal D}_1(t)$ and ${\mathcal
D}_2(t)$. The proof for the other cases can be dealt with
analogously. Define the two families
$$
X_{d_1,t}=\left\{\sum_{i=1}^5 y^{320}_i - \prod_i
y_i^{64}=0\right\},
$$
$$
X_{d_2,t}=\left\{\sum_{i=1}^5 y^{1280}_i - \prod_i
y_i^{256}=0\right\}.
$$

Let $A_1$ and $A_2$ be the following matrices:
$$
A_1:=\left(
\begin{array}{ccccc}
5 & 0 & 1 & 0 & 0 \\
0 & 0 & 4 & 1 & 0 \\
0 & 1 & 0 & 4 & 0 \\
0 & 4 & 0 & 0 & 1 \\
0 & 0 & 0 & 0 & 4
\end{array}
\right), \quad A_2:=\left(
\begin{array}{ccccc}
0 & 0 & 5 & 0 & 0 \\
5 & 0 & 0 & 1 & 0 \\
0 & 1 & 0 & 4 & 0 \\
0 & 1 & 0 & 0 & 4 \\
0 & 0 & 0 & 0 & 4
\end{array}
\right).
$$

Using Magma we found that the groups $H_{A_1}$ and $H_{A_2}$ are trivial - this
happens for all $12$ families.

We thus have
$$
q_{A_1} \circ \phi_{A_1} : X_{d_1, t} \rightarrow \overline{M}_{A_
1,t}$$
$$
(y_1: y_2: \ldots: y_5) \rightarrow  \left(\prod_i y_i^{64}:
y_1^{320}: \ldots : y_5^{320}\right)
$$
and, similarly,
$$
q_{A_2} \circ \phi_{A_2} : X_{d_2, t} \rightarrow \overline{M}_{A_
2,t}
$$
$$
(y_1: y_2: \ldots: y_5) \rightarrow  \left(\prod_i y_i^{256}:
y_1^{1280}: \ldots : y_5^{1280}\right)
$$

It is easy to check that $\overline{M}_{A_1,t} \cong
\overline{M}_{A_2,t} \cong \left\{ \sum_i u_i -tu_0, \,
u_0^5=u_1\ldots u_5\right\} \approx W_5(t)$.

Since $H_{A_1}$ and $H_{A_2}$ are trivial, ${\mathcal D}_1(t)$ and
${\mathcal D}_1(t)$ are birational since they are both birational to
$W_5(t)$.

\endproof

\subsection{Picard-Fuchs equations}\label{sec9}  When $X_A$ is a Calabi-Yau hypersurface,
the Hodge number $h^{2,1}(X_A)$ gives the number of independent
parameters of deformations of $X_A$. There exists a system of
partial differential equations, the so called GKZ-hypergeometric
system (see \cite{GKZ}), which yield Picard-Fuchs equations for the
variation of periods along families with central fiber $X_A$. When
$h^{2,1}(X_A)=1$, the Picard-Fuchs equation can also be found via a
generalization of the Griffiths-Dwork method for hypersurfaces in
weighted projective space: see, for instance, \cite{mor}.

\section{${\overline M}_A$ and the mirror family of Calabi-Yau hypersurfaces in weighted projective space}\label{sec5}

Assume $X_A$ is a Calabi-Yau manifold (as defined in Section 2) in
weighted projective space $W{\mathbb P}^{n-1}(q_1, \ldots, q_n)$,
where $q_i | Q$ and $Q=\sum_j q_j$.  Batyrev's mirror construction
(see, for instance, \cite{ph}) depends only on the polytope $\Delta$
associated to the toric variety $W{\mathbb P}^{n-1}(q_1, \ldots,
q_n)$ and not on the matrix $A$. As explained, for instance in
\cite{ph}, the Calabi-Yau varieties in the mirror family ${\mathcal
W} \rightarrow {\mathbb P}^1_x$ of a general section of the
anticanonical bundle ${\mathcal O}(Q)$, with $Q=\sum_j q_j$,  can be
represented as compactifications of complete intersections of the
affine hypersurfaces in $({\mathbb C}^*)^{n}$ given by
\begin{equation}
\label{okkei}
t_1 + \ldots + t_n=1, \qquad t_1^{q_1}\hdots t_n^{q_n}=x.
\end{equation}

Let $W_x$ be the fiber over the point $x \in {\mathbb P}^1$. By comparing \eqref{uzero} and \eqref{okkei}, the following holds.

\begin{proposition}
\label{mirro}
The compactification of ${\mathcal W}_1$ is given by
the equations \eqref{uzero} that define the Shioda quotient ${\overline M}_{{}^tA,1}$ for any matrix $A$.
\end{proposition}

\proof Let $A$ be a matrix as in Section \ref{sec1}. If we start from the family $F_{{}^tA, t}$,
the equations \eqref{uzero} become:
\begin{equation}
\sum_i u_i - t u_0=0, \qquad u^d_0 = u_1^{q_1}\ldots u_n^{q_n}.
\end{equation}

Since $\sum_j q_j=d$, the claim follows.
\endproof

\medskip

{\bf Acknowledgements.} The author wishes to express his gratitude
to Bert van Geemen for helpful remarks and useful suggestions.

 \end{document}